# Tokenizer Choice For LLM Training: Negligible or Crucial?

### v1

@@ Abstract
The recent success of Large Language Models (LLMs) has been predominantly driven by curating the training dataset composition, scaling of model architectures and dataset sizes and advancements in pretraining objectives, leaving tokenizer influence as a blind spot. Shedding light on this underexplored area, we conduct a comprehensive study on the influence of tokenizer choice on LLM downstream performance by training 24 mono- and multilingual LLMs at a 2.6 B parameter scale, ablating different tokenizer algorithms and parameterizations. Our studies highlight that the tokenizer choice can significantly impact the model’s downstream performance, training and inference costs. In particular, we find that the common tokenizer evaluation metrics fertility and parity are not always predictive of model downstream performance, rendering these metrics a questionable choice for tokenizer evaluation. Furthermore, we show that multilingual tokenizers trained on the five most frequent European languages require vocabulary size increases of factor three in comparison to English. While English-only tokenizers have been applied to the training of multi-lingual LLMs in the past, we find that this approach results in a severe downstream performance degradation and additional training costs of up to 68%, due to an inefficient tokenization vocabulary.
†Equal contribution.
@@ 1 Introduction
The quality of a tokenizer can be assessed intrinsically and extrinsically. An intrinsic evaluation solely addresses the characteristics of tokenizers and their generated output in isolation, whereas the extrinsic evaluation measures the impact of the tokenizer on a downstream component, e.g., the Large Language Model (LLM) model.
While many different tokenization approaches have been proposed, ranging from character-based to word-based methods, the potential impact of different tokenizers is neglected w.r.t. LLMs, especially in the context of multilingual LLMs. Recent work proposed by Petrov et al. (2023) demonstrates that carelessly designed tokenizers applied to the training of multilingual LLMs result in severe inequalities and limitations across languages. In fact, text passages translated into different languages resulted in tokenized sequences that differ in length up to a factor of 15, affecting inference costs and latency during inference. Furthermore, it is known that an essential property for effectively learning transformer-based LLMs is the learning of long-range dependencies Vaswani et al. (2017). Given a fixed sequence length, learning to relate words far apart in the input text is impossible for languages whose text is excessively fragmented by the tokenizer.
Despite the importance of tokenizers and the potentially severe impact of poorly performing tokenizers, there exists so far no extensive study that holistically investigates the intrinsic and extrinsic tokenizer performance in a monolingual and multilingual setting with a focus on decoder-only models, which represent the backbone of current LLMs.
Therefore, we address this gap and conduct an extensive study in which we measure the impact of the tokenizer on the model performance. In particular, we make the following contributions:
@@
  • Investigate whether a correlation between the intrinsic and the extrinsic tokenizer performance exists.
@@ 2.1 Tokenization Approaches
The most basic tokenization approach is to split sequences based on white spaces and consider each word as a token Bengio et al. (2000).
@@
Subword tokenization algorithms are data-driven tokenization approaches which can decompose words into subwords/multiple tokens and currently represent the established tokenization approach upon which LLMs rely Kudo and Richardson (2018); Petrov et al. (2023). Because subword tokenizers decompose words into subwords, they can process out-of-vocabulary words by merging subwords from the vocabulary Kudo and Richardson (2018). Examples of popular subword tokenizers are WordPiece Schuster and Nakajima (2012), BPE Gage (1994); Sennrich et al. (2015), Byte-Level BPE (BBPE) Wang et al. (2020), and Unigram Kudo (2018).
@@ 2.2 Tokenizers in Transformers Models
An overview of current mono- and multilingual LLMs is provided in Lin et al. (2022); Shliazhko et al. (2022); Scao et al. (2022). As part of their work, Shliazhko et al. (2022) ablated different tokenizer pre-processing approaches while keeping the tokenizer algorithm, the vocabulary size and the employed implementation fixed. In none of the other major LLMs works the extrinsic tokenizer performance has been studied.
@@ 3.3 Models
To measure the impact of our trained tokenizers on the model downstream performance, we trained one model for each tokenizer. In particular, for each of our 24 trained tokenizers, we trained a 2.6B transformer-based decoder-only model on up to 52B tokens following the scaling law proposed by Hoffmann et al. (2022a). Additionally, serving as baselines, we trained a monolingual and a multilingual model using the pre-trained GPT-2 tokenizer. All models have been trained based on the causal language modeling training objective.
@@ 3.4 Evaluation
Parity Petrov et al. (2023), which has been recently proposed, assesses how fairly a tokenizer treats equivalent sentences in different languages. A tokenizer $T$ achieves parity for language $A$ with respect to language $B$ if $\frac{|T(s_{A})|}{|T(s_{B})|}\approx 1$, where $s_{A}$ and $s_{B}$ denote the sets of all sentences in the corpora of language $A$ and $B$, respectively. We use the FLORES-200 Goyal et al. (2022) parallel corpus, consisting of the same sentences human-translated into 200 languages. We calculate the parity values for each tokenizer and the four non-English languages with respect to English (see Fig. 2 for an overview).
The extrinsic evaluation aims to explicitly assess the impact of a tokenizer on the model’s downstream performance. We selected a comprehensive set of downstream tasks covering commonsense reasoning, reading comprehension, translation, and math tasks to measure the downstream performance. For an exhaustive lists of all tasks, we refer the reader to Table 11 in Appendix E.
Additionally, we computed the impact of a tokenizer on the average computational costs of a given model per word during training and inference. The computational costs during training for one step including the forward and the backward pass can be estimated by
@@
given a model with batch size $B$, sequence length $s$, $l$ layers, hidden size $h$ and vocabulary size $V$ Narayanan et al. (2021). The costs per token can be derived by $C_{\text{token}}=C/Bs$ and the average costs per word by $C_{\text{word}}=C_{\text{token}}\cdot\text{fertility}$.
@@ 4 Intrinsic Tokenizer Evaluation
In our intrinsic evaluation, we first compare the fertility and parity of the trained tokenizers (Section 4.1) and the overlap of their vocabularies (Section 4.2).
@@ 4.1 Fertility & Parity
Figure 1: Comparison of fertility scores between monolingual (English) tokenizer and multilingual tokenizers applied to (a) English and (b) non-English, multi-lingual documents.
Applying the described fertility and parity evaluation to the mono-/multilingual tokenizers, our analysis highlights the following two major aspects, as visible in Fig. 1 and Fig. 2.
@@
Secondly, with increasing vocabulary size, fertility and parity reduce in all cases, which can be explained by the tokenizer requiring fewer sub-word tokens when tokenizing text given a larger vocabulary. However, it can be observed that for monolingual English tokenizers, the fertility is less dependent on the vocabulary when tokenizing English documents, implying that 33k might be a sufficiently large vocabulary. This is especially interesting considering the fact that the LLaMA/LLaMA2 models have been trained with a vocabulary size of 33k Touvron et al. ; Touvron et al. (2023).
@@ 4.2 Vocabulary Overlap
To analyze the tokenizer similarity, we calculated the vocabulary overlap. Particularly, we assess Huggingface’s and SentencePiece’s BPE implementations. Their vocabulary overlap is depicted in Table 1.
@@
(b) English documents
(c) German documents
@@ 5 Extrinsic Tokenizer Evaluation
In the following, we describe the results of our extrinsic evaluation of tokenizers. Section 5.1 describes the experimental setup, Section 5.2 presents the downstream performance of the trained models based on the investigated tokenizers, and Section 5.3 analyses the computational costs associated with each tokenizer when employed in a specific model.
@@ 5.1 Experimental Setup
To assess the impact of the tokenizers on the model downstream performance, we trained a decoder-only transformer model of size 2.6 B for each tokenizer. We trained our models for 52.6 B tokens following the scaling law proposed by Hoffman et al. Hoffmann et al. (2022b) based on the causal language modeling training objective. The hyper-parameters are described in Table 8 in the Appendix C.
We evaluated our models in zero-shot settings on a wide range of tasks, covering commonsense reasoning (HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2020), ARC easy and challenge Clark et al. (2018)), world knowledge (TriviaQA Joshi et al. (2017)), reading comprehension (LAMBADA Paperno et al. (2016) and their machine translated versions for German, Spanish, French, and Italian, RACE Lai et al. (2017), BoolQ Clark et al. (2019)), math (MATH Hendrycks et al. (2021)), multilingual natural language inference (XNLI Conneau et al. (2018)).
@@ 5.2 Downstream Performance
We split our analysis of the downstream performance into several parts. First, we discuss the overall results obtained for the investigated tokenizers, followed by presenting the impact of the tokenizer library (Section 5.2.1), the impact of the tokenizer algorithm (Section 5.2.2), and the impact of the vocabulary size (Section  5.2.3).
(a) Monolingual tokenizers.
(b) Multilingual tokenizers.
Figure 4: Top-1/-2/-3 count for mono-/multilingual tokenizers across mono-/multilingual tasks the tasks were evaluated on the accuracy metric.
Model DE FR IT ES EN
EN BPE-HF - - - - 45.44
BPE-SP - - - - 46.31
UNI-SP - - - - 46.01
MULTI BPE-HF 36.72 36.55 35.34 42.28 44.45
BPE-SP 36.96 36.95 36.12 41.96 44.92
UNI-SP 37.38 37.05 35.70 42.56 44.21
Table 2: Impact of the tokenizer algorithm and the tokenizer training library, scores are averaged over the vocabulary sizes.
Vocab DE FR IT ES EN
EN 32 - - - - 46.34
50 - - - - 46.30
82 - - - - 45.29
100 - - - - 45.88
MULTI 32 36.72 36.22 35.20 41.76 44.71
50 36.57 36.34 34.60 42.22 43.91
82 37.40 37.22 36.21 42.30 44.75
100 37.17 37.45 36.19 42.64 44.55
Table 3: Impact of the vocabulary size on the downstream accuracy. The accuracy scores are averaged over the libraries and tokenizer algorithms
To provide a holistic performance overview, we assess the impact of the tokenizers on the model performance based on various tasks and analyze the results from different perspectives instead of providing solely a single aggregated metric.
This holistic evaluation is presented in Fig. 4, which expresses how often a tokenizer is among the top-1/-2/-3 performing tokenizers, Fig. 7, which shows the distribution of the accuracy scores across tasks, and in Table 9 which summarizes the average performances across all tasks. We dissect the results from different perspectives in the following subsections.
@@
Among the monolingual tokenizers, there can be significant performance differences. For instance, Fig. 3(a) depicts that the GPT-2 tokenizer is overall most often across the top-$k$ (with $k\in\{1,2,3\}$) performing tokenizers, whereas the EN-BPE-HF-32 tokenizer is never the best-performing tokenizer. Though the count-based metric already provides a first overview of the model performance, it does not illustrate the difference in performance.
Therefore, we also looked at the accuracy distribution across all tasks that employ accuracy as the evaluation metric. Fig. 6(a) demonstrates that the EN-BPE-SP-32 tokenizer, on average, is across all tasks among the best-performing tokenizers and has the highest median value. At the same time, it also expresses a large variance across tasks compared to other tokenizers such as EN-BPE-HF-32, implying that for a set of tasks, the tokenizer performs very well, while on other tasks, it performs comparably poorly.
Aggregated metrics provide a good impression of the overall performance but do not demonstrate potentially large performance differences for single tasks. Therefore, we listed in Table 11 the performance difference across all tasks between the best-performing tokenizer and worst-performing tokenizer on this task. The performance difference is huge for various tasks, which is especially interesting since only the tokenizer has been changed while the model configuration is the same.
Overall, the GPT-2 tokenizer and the EN-BPE-SP-32 tokenizer are strong tokenizers for monolingual English models. Interestingly, SentencePiece’s implementation of BPE with a vocabulary size of 32k has been used for LLaMA2. Our results provide empirical and transparent evidence that both tokenizers/tokenizer settings are reasonable choices for training monolingual English models.
@@
Our multilingual tokenizer experiments further support that different tokenizers can express large performance differences. Based on our count-based metric (Figure 3(b)), the differences across tokenizers are even larger than for monolingual tokenizers. While Multi-UNI-SP-100 and Multi-UNI-SP-82 are most often among the strongest tokenizers, Multi-BPE-HF-32 is only a single time across the top-performing tokenizers.
The accuracy distributions depicted in Figure 6(b) further show that both the tokenizers mentioned above are among the best-performing tokenizers and that specific tokenizers such as the already mentioned tokenizer perform worse. Figure 6(b) additionally demonstrates that the GPT-2 tokenizer performs poorly. Compared to monolingual models, we can observe that all tokenizers express a similar variance across tasks.
In reference to Table  11, the analysis of single tasks reveals that for multilingual tokenizers, the performance difference between tasks can be huge.
Multi-UNI-SP-100 and Multi-UNI-SP-82 are excellent choices, while employing a pre-trained GPT-2 tokenizer should be omitted to train multilingual models.

### v4

@@ Abstract
The recent success of Large Language Models (LLMs) has been predominantly driven by curating the training dataset composition, scaling of model architectures and dataset sizes and advancements in pretraining objectives, leaving tokenizer influence as a blind spot. Shedding light on this underexplored area, we conduct a comprehensive study on the influence of tokenizer choice on LLM downstream performance by training 24 mono- and multilingual LLMs at a 2.6 B parameter scale, ablating different tokenizer algorithms and parameterizations. Our studies highlight that the tokenizer choice can significantly impact the model’s downstream performance and training costs. In particular, we find that the common tokenizer evaluation metrics fertility and parity are not always predictive of model downstream performance, rendering these metrics a questionable proxy for the model’s downstream performance. Furthermore, we show that multilingual tokenizers trained on the five most frequent European languages require vocabulary size increases of factor three in comparison to English. While English-centric tokenizers have been applied to the training of multi-lingual LLMs in the past, we find that this approach results in a severe downstream performance degradation and additional training costs of up to 68%, due to an inefficient tokenization vocabulary.
Tokenizer Choice For LLM Training: Negligible or Crucial?
Mehdi Ali1,2 ${}^{\dagger}$, Michael Fromm1,2 ${}^{\dagger}$, Klaudia Thellmann3 ${}^{\dagger}$ Richard Rutmann1,2, Max Lübbering1,2, Johannes Leveling1, Katrin Klug1, Jan Ebert4, Niclas Doll1, Jasper Schulze Buschhoff1, Charvi Jain1,2, Alexander Arno Weber1,2, Lena Jurkschat3, Hammam Abdelwahab1 Chelsea John4, Pedro Ortiz Suarez5, Malte Ostendorff5 Samuel Weinbach6, Rafet Sifa1, Stefan Kesselheim4, Nicolas Flores-Herr1 1Fraunhofer IAIS, 2Lamarr Institute, 3TU-Dresden, 4FZ Jülich, 5DFKI, 6Aleph Alpha thanks: †Equal contribution.
@@ 1 Introduction
The quality of a tokenizer can be assessed intrinsically and extrinsically. An intrinsic evaluation solely addresses the characteristics of tokenizers and their generated output in isolation, whereas the extrinsic evaluation measures the impact of the tokenizer on a downstream component, e.g., the Large Language Model (LLM).
While many different tokenization approaches have been proposed, ranging from character-based to word-based methods, the potential impact of different tokenizers is underexplored w.r.t. LLMs, especially in the context of multilingual LLMs. Recent work proposed by Petrov et al. (2023) demonstrates that carelessly designed tokenizers applied to the training of multilingual LLMs result in severe inequalities and limitations across languages. Text passages translated into different languages resulted in tokenized sequences that differ in length up to a factor of 15, affecting inference costs and latency during inference. Furthermore, it is known that the learning of long-range dependencies Vaswani et al. (2017), is an essential property for effectively learning transformer-based LLMs. Given a fixed sequence length, learning to relate words far apart in the input text is impossible for languages whose text is excessively fragmented by the tokenizer.
Despite the importance of tokenizers and the potentially severe impact of poorly performing tokenizers, there exists no extensive study so far that holistically investigates the intrinsic and extrinsic tokenizer performance in a monolingual and multilingual setting with a focus on decoder-only models, which represent the backbone of current LLMs.
In this work, we address this gap and conduct an extensive study in which we measure the impact of the tokenizer on the model performance. In particular, we make the following contributions:
@@
  • We investigate whether a correlation between the intrinsic and the extrinsic tokenizer performance exists.
@@ 2.1 Tokenization Approaches
The most basic tokenization approach is the splitting of sequences based on white spaces and considering each word as a token Bengio et al. (2000).
@@
This class of algorithms subsumes all data-driven tokenization approaches which can decompose words into subwords/multiple tokens and currently represent the established tokenization approach upon which LLMs rely Kudo and Richardson (2018); Petrov et al. (2023). Because subword tokenizers decompose words into subwords, they can process out-of-vocabulary words by merging subwords from the vocabulary Kudo and Richardson (2018). Examples of popular subword tokenizers are WordPiece Schuster and Nakajima (2012), BPE Gage (1994); Sennrich et al. (2015), Byte-Level BPE (BBPE) Wang et al. (2020), and Unigram Kudo (2018).
@@ 2.2 Tokenizers in Transformers Models
An overview of current mono- and multilingual LLMs is provided in Lin et al. (2022); Shliazhko et al. (2022); Scao et al. (2022). Stollenwerk (2023) evaluated the intrinsic metrics of the GPT-SW3 Ekgren et al. (2023) tokenizer that focused on the Nordic languages. As part of their work, Shliazhko et al. (2022) ablated different tokenizer pre-processing approaches while keeping the tokenizer algorithm, the vocabulary size, and the employed implementation fixed. In none of the other major LLM publications, the extrinsic tokenizer performance has been studied.
@@ 3.3 Models
To measure the impact of our trained tokenizers on the model downstream performance, we trained one model for each tokenizer. In particular, for each of our 24 trained tokenizers, we trained a 2.6B transformer-based decoder-only model on up to 52B tokens following the scaling law proposed by Hoffmann et al. (2022a). Additionally, serving as baselines, we trained a monolingual and a multilingual model using the pre-trained GPT-2 tokenizer Radford et al. (2018). All models have been trained based on the causal language modeling training objective.
@@ 3.4 Evaluation
Parity Petrov et al. (2023), which has been recently proposed, assesses how fairly a tokenizer treats equivalent sentences in different languages. A tokenizer $T$ achieves parity for language $A$ with respect to language $B$ if $\frac{|T(s_{A})|}{|T(s_{B})|}\approx 1$, where $s_{A}$ and $s_{B}$ denote the sets of all sentences in the corpora of languages $A$ and $B$, respectively, and the ratio $\frac{|T(s_{A})|}{|T(s_{B})|}$ is defined as premium. We use the FLORES-200 Goyal et al. (2022) parallel corpus, consisting of the same sentences human-translated into 200 languages. We calculate the parity values for each tokenizer and the four non-English languages with respect to English (see Fig. 2 for an overview).
The extrinsic evaluation aims to explicitly assess the impact of a tokenizer on the model’s downstream performance. We selected a comprehensive set of downstream tasks (see Section 5.1) to measure the downstream performance.
Additionally, we computed the impact of a tokenizer on the average computational costs of a given model per word during training. The computational costs during training for one step including the forward and the backward pass can be estimated by
@@
given a model with batch size $B$, sequence length $s$, $l$ layers, hidden size $h$ and vocabulary size $V$ Narayanan et al. (2021). The costs per token can be derived by $C_{\text{token}}=C/Bs$ and the average costs per word by $C_{\text{word}}=C_{\text{token}}\times\text{fertility}$. The Results are discussed in Section 5.3.
@@ 4 Intrinsic Tokenizer Evaluation
In our intrinsic evaluation, we first compare the fertility and parity of the trained tokenizers (Section 4.1) and subsequently the overlap of their vocabularies (Section 4.2).
@@ 4.1 Fertility & Parity
Figure 1: Comparison of fertility scores between mono- and multilingual tokenizers applied to (a) Non-English, multilingual documents and (b) English documents.
Applying the described fertility and parity evaluation to the mono-/multilingual tokenizers, our analysis highlights the following two major aspects, as visualized in Fig. 1 and Fig. 2.
@@
Secondly, with increasing vocabulary size, fertility and parity reduce in all cases, which can be explained by the tokenizer requiring fewer sub-word tokens when tokenizing text given a larger vocabulary. However, it can be observed that for monolingual English tokenizers, the fertility is less dependent on the vocabulary when tokenizing English documents, implying that 33k might be a sufficiently large vocabulary.
@@ 4.2 Vocabulary Overlap
To analyze the tokenizer similarity, we calculated the vocabulary overlap. Particularly, we assess Huggingface’s and SentencePiece’s BPE implementations, as depicted in Table 1.
@@
(b) German documents
(c) English documents
@@ 5 Extrinsic Tokenizer Evaluation
In the following, we describe the results of our extrinsic evaluation of tokenizers. Section 5.1 describes the experimental setup, Section 5.2 presents the downstream performance of the trained models based on the investigated tokenizers, and Section 5.3 analyzes the computational costs associated with each tokenizer when employed in a specific model.
@@ 5.1 Experimental Setup
To assess the impact of the tokenizers on the model downstream performance, we trained a decoder-only transformer model of size 2.6 B for each tokenizer. We trained our models for 52.6 B tokens following the scaling laws proposed by Hoffmann et al. (2022b), based on the causal language modeling training objective. The hyper-parameters are described in Table 10 in the Appendix C. We evaluated our models in zero-shot settings on a wide range of mono- and multilingual tasks:
  • Natural language inference: XNLI Conneau et al. (2018), MNLI  Williams et al. (2018), RTE Wang et al. (2018), WNLI Levesque et al. (2012), CB De Marneffe et al. (2019)
  • Question answering: X-CSQA Goodman (2001), XStoryCloze Lin et al. (2022), PubMedQA Jin et al. (2019)
  • Reading comprehension: BoolQ Clark et al. (2019)), LAMBADA Paperno et al. (2016), RACE Lai et al. (2017), MRPC Dolan and Brockett (2005).
  • Commonsense reasoning: HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2020), ARC Clark et al. (2018), XCOPA Ponti et al. (2020), XCDOAH Goodman (2001), WSC Levesque et al. (2012), COPA Roemmele et al. (2011)
  • Classification: PAWS-X Yang et al. (2019), GNAD10 Schabus et al. (2017), SST Socher et al. (2013), WIC Pilehvar and Camacho-Collados (2019), PIQA Bisk et al. (2020)
Table 2 provides an overview of the number of tasks for each category and language.
Task EN DE FR ES IT
NLI 6 1 1 1 0
QA 3 2 2 3 2
RC 3 1 1 1 1
CR 7 0 1 0 1
CL 3 1 0 1 0
22 5 4 6 4
Table 2: Overview of the number of evaluation tasks for each language and the categories of Natural language inference (NLI), Reading comprehension (RC), Question answering (QA), Commonsense reasoning (CR) and Classification (CL).
@@ 5.2 Downstream Performance
We split our analysis of the downstream performance into several parts.
First, we discuss the overall results obtained for the investigated tokenizers, followed by presenting the impact of the tokenizer library (Section 5.2.1), the impact of the tokenizer algorithm (Section 5.2.2), and the impact of the vocabulary size (Section 5.2.3).
We present both, aggregated results across all tasks (Table 3) and results for selected single tasks (Table 4). For the average performance across all tasks presented in Table 3, we computed weighted average to take into account the different number of tasks per language. In particular, we computed for each language the mean across all tasks, and then computed the mean over all language-means.
Model EN MULTI
GPT-2-50 50.36 39.41
BPE-HF-33 49.13 40.52
BPE-HF-50 49.51 40.47
BPE-HF-82 48.71 40.24
BPE-HF-100 49.54 40.48
BPE-SP-33 50.81 40.28
BPE-SP-50 49.81 40.49
BPE-SP-82 48.99 41.21
BPE-SP-100 49.46 41.44
UNI-SP-33 50.28 40.30
UNI-SP-50 49.90 40.48
UNI-SP-82 49.65 41.20
UNI-SP-100 50.21 40.74
Table 3: Average accuracy of monolingual and multilingual tokenizers across all downstream tasks. Due to varying number of tasks per language, multi-lingual accuracies have been adjusted to each language contributing equally to the average.
Task Min Max Rand.
EN ARC-Easy 0.50 0.59 0.20
HellaSwag 0.34 0.41 0.25
MRPC 0.54 0.69 0.50
PIQA 0.67 0.72 0.50
MULTI XNLI FR 0.37 0.49 0.33
XNLI EN 0.49 0.52 0.33
X-CODAH ES 0.28 0.43 0.25
10kGNAD 0.15 0.43 0.11
Table 4: Worst- and best-performing tokenizer for selected tasks and the random performance on this task.
MULTI MONO
Vocabulary DE FR IT ES EN AVG EN
33 36.75 36.66 39.30 41.76 47.37 40.37 49.55
50 36.12 37.07 38.94 42.22 46.71 40.21 49.90
82 36.50 37.83 39.97 42.30 47.80 40.88 49.12
100 35.92 38.07 40.13 42.64 47.67 40.89 49.74
Algorithm and Library DE FR IT ES EN AVG EN
BPE-HF 35.69 37.31 39.37 42.28 47.48 40.43 48.98
BPE-SP 37.13 37.45 40.04 41.96 47.68 40.85 49.77
UNI-SP 36.51 37.66 39.57 42.56 47.10 40.68 50.01
Table 5: Impact of the vocabulary size (upper), and tokenizer algorithm and library (lower), on the downstream performance. The accuracy scores are either averaged over the libraries and tokenizer algorithms (upper) or the different vocabulary sizes (lower).
@@
Table 3 demonstrates that the BPE-SP-33 tokenizer, on average, is the best-performing tokenizer, followed by the GPT-2 tokenizer. Interestingly, SentencePiece’s implementation of BPE with a vocabulary size of 33k has been used for LLaMA2 Touvron et al. (2023). Aggregated metrics provide a reasonable overview of the overall performance. However, it does not express potentially large performance differences across tasks. Therefore, we listed in Table 4 the obtained results for a list of selected tasks obtained by the best and worst performing tokenizer on this task. The results illustrate that the performance difference can be huge. For instance, for ARC-Easy, a commonsense reasoning task, the gap between the best and worst tokenizer is 9%.
@@
Table 3 shows that the BPE-SP-100 tokenizer is the best-performing tokenizer followed by the BPE-SP-82 tokenizer. Furthermore, Table 3 demonstrates that the GPT-2 tokenizer performs poorly, implying that using a pre-trained GPT-2 tokenizer to pre-train and fine-tune multilingual models should be omitted. The analysis of selected tasks ( 4) reveals that for multilingual tokenizers, the performance difference between tasks can be huge.